The product development cost of Canon CI-700 is as follows. Suppose production volume is 5,000 per...

  1. The product development cost of Canon CI-700 is as follows. Suppose production volume is 5,000 per year, what is annual profit? How many years does it take to break-even? (10pt)

Setup cost

$3 million

Ramp-up cost

$2 million

Marketing and support cost

$1 million/year

Unit production cost

$400/unit

Unit price

$800/unit
Answer #1

Fixed costs here are setup costs, Ramp up costs, marketing costs which all add upto 6 million dollar per year.

Variable costs here are unit production cost =400×5000=2 million dollar per year

Now selling price for all units= 800×5000= 4 million dollar

Thus,

Now since selling price is 4milliin dollars and total costs are 8 million dollars, it will take 2 years forno profit no loss situation that is in other words called breakeven.